Why the speed of sound in water is more than mercury, though the density of mercury is more than water?

1
Expert's answer
2021-12-20T10:30:03-0500

The speed of sound in liquids depends not only on the density but also on the bulk modulus "K" . So,


"v=\sqrt{\frac{K}{\rho}}"


"v_{water}=\sqrt{\frac{K}{\rho}}=\sqrt{\frac{2.2\cdot10^9}{1000}}=1483\ (m/s)"


"v_{Hg}=\sqrt{\frac{K}{\rho}}=\sqrt{\frac{25\cdot10^9}{13520}}=1360\ (m/s)"